Five times a number k minus 9 is two thirds of a number j. Solve for j.
Answer by nerdybill(7384)

Five times a number k minus 9 is two thirds of a number j. Solve for j.
.
Translating the "English" to math:
5k-9 = (2/3)j
Now, we solve for j. Start, by multiplying both sides by 3:
15k-27 = 2j
Now, divide both sides by 2:
(15/2)k-(27/2) = j